OpenStudy (megannicole51):

Set up a double integral that corresponds to the mass of the semi disk (give me a min to post the formula)

OpenStudy (megannicole51):

\[D=(x,y) : 0 \le y \le \sqrt{4-x ^{2}}\] if the density at each of its points is given by the function \[\delta (x,y)= x ^{2}y\]

OpenStudy (megannicole51):

\[\int\limits_{-2}^{2} \int\limits_{0}^{\sqrt{4-x ^{2}}} x ^{2}y* dy dx\] this is the answer

OpenStudy (megannicole51):

@hartnn

OpenStudy (megannicole51):

anybodyyyyyyy

OpenStudy (agent0smith):

Well this one I kinda get, mass is the integral of density over volume.

OpenStudy (agent0smith):

The y limits come from this \[\Large 0 \le y \le \sqrt{4-x ^{2}}\]so that gives \[\Large \int\limits\limits_{0}^{\sqrt{4-x ^{2}}} x ^{2}y* dy\](integrating the density over the y limits)

OpenStudy (megannicole51):

the next question is find the mass of the above semidisk

OpenStudy (megannicole51):

yeah but how did they know where the endpoint to the integral go for a and b?

OpenStudy (agent0smith):

And the x-limits come from the \[\Large \sqrt{4-x ^{2}} \]since you can only get real numbers for y if -2<x<2 right?

OpenStudy (megannicole51):

yeah but is that for every case?

OpenStudy (agent0smith):

Every case...?

OpenStudy (megannicole51):

and how do u find the mass now?

OpenStudy (megannicole51):

every time i get these kinds of problems?

OpenStudy (agent0smith):

In this case it's because of that sqrt(4-x^2)... it's only real numbers if x is between -2 and 2, since square roots of negative numbers are imaginary. Okay the mass means we need to integrate this \[\Large \int\limits\limits_{-2}^{2} \int\limits\limits_{0}^{\sqrt{4-x ^{2}}} x ^{2}y* dy dx\]

OpenStudy (agent0smith):

It might make it easier to change the order of integration (you can do that with double integrals), do dx first then dy. \[\Large \int\limits\limits_{0}^{\sqrt{4-x ^{2}}}\int\limits\limits_{-2}^{2} x ^{2}y* dx dy \]

hartnn (hartnn):

you know you can't do that ^, right ?

OpenStudy (agent0smith):

Wait... you can't? I thought it was valid...

hartnn (hartnn):

i mean you can't change the order like that..

hartnn (hartnn):

and why do u even need to change the order ?

hartnn (hartnn):

its a really easy double integral to solve, integrate w.r.t y first, treating x as constant

OpenStudy (megannicole51):

i took the integral from 0 to the sqrt 4-x^2 and integrated it which came to x^3*y/3 and then dont i just solve it from 0 to the sqrt?

OpenStudy (megannicole51):

and then once i get that answer integrate it from -2 to 2?

OpenStudy (agent0smith):

\[\Large \int\limits\limits_{0}^{\sqrt{4-x ^{2}}} x ^{2}y* dy = \left[ \frac{ 1 }{ 2}x^2 y^2 \right]_{0}^{\sqrt{4-x^2}}\]oh right i was thinking the square root makes things awkward but it doesn't

OpenStudy (agent0smith):

Meg you integrated wrt x not y

OpenStudy (megannicole51):

whats the difference? ive nvr been taught that...well tht i know of

OpenStudy (agent0smith):

\[\Large \left[ \frac{ 1 }{ 2}x^2 y^2 \right]_{0}^{\sqrt{4-x^2}} = \frac{ 1 }{ 2}x^2(4-x^2)\]

OpenStudy (megannicole51):

wait idk how to integrate that

OpenStudy (agent0smith):

dy means you integrate y, not x. Like y integrates to y^2/2, same way x integrates to x^2/2

OpenStudy (anonymous):

well x is from -2 to 2, and y is from 0 to sqrt(4-x^2), you have the limit of integration.

OpenStudy (megannicole51):

how did u come up with that integration?

OpenStudy (agent0smith):

Treat x as a constant. increase power on y, divide by new power\[\Large \int\limits\limits\limits_{0}^{\sqrt{4-x ^{2}}} x ^{2}y* dy = \left[ \frac{ 1 }{ 2}x^2 y^2 \right]_{0}^{\sqrt{4-x^2}}\]

OpenStudy (megannicole51):

still dont get it

OpenStudy (megannicole51):

:/ sorry

OpenStudy (megannicole51):

oh wait i see it

hartnn (hartnn):